A company had total revenues of $129 million, operating profit margin of 19%, and depreciation and amortization expense of $15 million over the trailing twelve months. The company currently has $39 million in total debt and $13 million in cash and cash equivalents. If the company's market capitalization (market value of its equity) is $543 million, what is its EV/EBITDA ratio? Round to one decimal place.
Answer #1

EBITDA=Revenues*Operating Profit Margin+Depreciation=129*19%+15=39.51 million

EV=Equity+Debt-Cash=543+39-13=569

EV/EBITDA=569/39.51=14.40141736